At an amusement park there is a ride in which cylindrically shaped chambers spin around a central axis. People sit in seats facing the axis, their backs against the outer wall. At one instant the outer wall moves at a speed of 2.79 m/s, and an 59.5-kg person feels a 191-N force pressing against his back. What is the radius of a chamber?
Answer #1

using the formula for centripetal force F= mv2/r

we get r =mv2/F= 59.5 *2.79*2.79/191 =2.425 m ans